OpenStudy (anonymous):

What is the value of x, when 13 = square root of(4x + 2) ?

OpenStudy (amistre64):

13 = sqrt(4x+2) 169 = 4x+2 167 = 4x x = 167/4 = abt 42

OpenStudy (amistre64):

41' 3/4